On January 2, 2018, Gold Star Leasing Company leases equipment to Brick Co. with 5 equal annual payments of $160,000 each, payable beginning January 2, 2018. Brick Co. agrees to guarantee the $150,000 residual value of the asset at the end of the lease term. The expected value of the residual value is $50,000. Brick's incremental borrowing rate is 10%, however it knows that Gold Star's implicit interest rate is 8%. What journal entry would Brick Co. make at January 2, 2018 to record the lease?

PV Annuity Due           PV Ordinary Annuity               PV Single Sum

            8%, 5 periods                    4.31213                          3.99271                                 .68508

            10%, 5 periods                  4.16986                          3.79079                                 .62092

a. Right-of-Use Asset 598,449

Lease Liability 598,449

b. Right-of-Use Asset 758,449

Cash 160,000

Lease Liability 598,449

c. Right-of-Use Asset 689,940

Cash 160,000

Lease Liability 529,940

d. Right-of-Use Asset 707,342

Cash 160,000

Lease Liability 547,342
Answer #1
PV of annuity payments 689941 =160000*4.31213
PV of residual value 68508 =(150000-50000)*0.68508
Right-of-Use Asset 758449 =689941+68508
Right-of-Use Asset 758449
        Cash 160000
        Lease Liability 598449
Option B is correct